# Attacking LLM Watermarks by Exploiting Their Strengths

### v1

@@ 6.3 Defending Detection with Differential Privacy
To better protect public detection APIs, we propose an effective defense using ideas from differential privacy (DP) [DR+14] to counteract spoofing attacks. DP adds random noise to function results evaluated on private dataset. Such that the results from neighboring datasets are indistinguishable. We apply this idea by adding Gaussian noise to the distance score in the watermark detection, and make the detection $(\epsilon,\delta)$-DP [DR+14]. Such that the attackers cannot tell the difference between two queries by replacing a single token in the content, which increases the hardness of launching the attacks. In the following, we evaluate the utility of the DP defense and its performance in mitigating spoofing attacks.
(a) Spoofing attack success rate and detection accuracy.
(b) Z-scores with/without DP and under multiple queries.
Figure 4: Evaluation of DP detection on KGW watermark and Llama-2-7B model. (a). Spoofing attack success rate (ASR) and detection accuracy (ACC) without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP under different query numbers (marked as 1×1$\times$1 ×Q, 20×20$\times$20 ×Q, and 200×200$\times$200 ×Q). We use the best $\sigma=4$ from (a).
[…]
@@ 6.4 Discussion
Guideline #3 DP techniques can help to mitigate spoofing attacks exploiting detection APIs. However, we also suggest that detection services should aim to verify users’ identities, identify malicious behavior, and limit query rates from potential attackers given the feasibility of both spoofing and removal attacks.
@@
Watermark detection APIs, which are typically available to the public, can be helpful to aid users in differentiating between AI-generated and human-created materials. However, these APIs can be exploited by malicious users to gradually remove watermarks or launch spoofing attacks. We propose a defense by employing ideas from differential privacy, which can make it significantly more difficult for attackers to launch spoofing attacks. However, this method is less effective against watermark-removal attacks that exploit the detection API because attackers’ actions will be close to random sampling, which may still prove effective when removing watermarks in practice. We thus leave developing a more powerful defense mechanism against watermark-removal attacks exploiting detection API as future work, and as a simple defense, recommend that companies providing detection services should detect and curb malicious behavior by limiting query rates from potential attackers and verifying the identity of the users to protect against Sybil attacks.
7 Conclusion & Discussion
@@
In this work, we reveal and evaluate new attack vectors that exploit the common properties of LLM watermarks. In particular, while these properties may enhance robustness, output quality, and public detection ease, they also allow malicious actors to launch attacks that can easily remove the watermark or damage the model’s reputation. Based on the theoretical and empirical analysis of our attacks, we suggest guidelines for designing and deploying LLM watermarks along with possible defenses to establish more reliable watermark embedding and detection systems.
@@
Our work studies the security implications of common LLM watermarking design choices. By developing realistic attacks and defenses and a simple set of guidelines for watermarking in practice, we aim for the work to serve as a resource for the development of secure LLM watermarking systems. Of course, we note that by outlining such attacks, there is a risk that our work may in fact increase the prevalence of watermark removal or spoofing attacks performed in practice. We believe that this is nonetheless an important step towards educating the community about potential risks in watermarking systems and ultimately creating more effective defenses for secure LLM watermarking.
@@
More generally, our work shows that a number of trade-offs exist in LLM watermarking (e.g., between quality, robustness, detection effectiveness, and susceptibility to removal or spoofing attacks). The guidelines we propose provide rough proposals for considering these trade-offs, but we note that how to best navigate each trade-off will likely depend on the application at hand. Considering additional strategies to navigate this space for specific LLM watermarking applications is an important direction for future study.
Appendix A Watermarking Schemes
@@
In this section, we introduce the three watermarking schemes we evaluate in the paper—KGW [KGW+23a], Unigram [ZALW24], and Exp [KTHL23]. We also introduce the perplexity, a metric to evaluate the sentence quality.
@@
KGW. In the KGW watermarking scheme, when generating the current token $\textbf{x}_{t+1}$, all the tokens in the vocabulary are pseudorandomly shuffled and split into two lists—the green list and the red list. The random seed used to determine the green and red lists is computed by a watermark secret key $sk$ the prior token $\textbf{x}_{t}$ using Pseudorandom functions (PRFs):
$\textsc{seed}=F_{sk}(\textbf{x}_{t})$
@@
Then, the seed is used to split the vocabulary into the green and red lists of tokens, with $\gamma$ portion of tokens in the green list:
@@
Then, KGW generates a binary watermark mask vector for the current token prediction, which has the same size as the vocabulary. All the tokens in the green list $L_{\text{green}}$ have a value 111 in the mask, and all the tokens in the red list have a value 0 in the mask:
@@
To embed the watermark, KGW adds a constant to the logits of the LLM’s prediction for token $\textbf{x}_{t+1}$:
@@
where $g$ is the number of tokens in the green list, $l$ is the total number of tokens in the input token sequence, and $\gamma$ is the portion of the vocabulary tokens in the green list. Similar to the watermark embedding, the green and red lists for each token position are determined by the watermark secret key and the token prior to the current token in the input token sequence.
Unigram. Similar to KGW, Unigram also splits the vocabulary into green and red lists and prioritizes the tokens in the green list by adding a constant to the logits before computing the softmax. The difference is that Unigram uses global red and green lists instead of computing the green and red lists for each token. That is, the seed to shuffle the list is only determined by the watermark secret key and generated by a Pseudo-Random Generator (PRG):
@@
The watermark embedding and detection procedures are the same as KGW: Unigram first computes the watermark mask:
@@
where $g$ is the number of tokens in the green list, $l$ is the total number of tokens in the input token sequence, and $\gamma$ is the portion of the vocabulary tokens in the green list. According to the analysis in [ZALW24] and also consistent with our results in Sec. 4.1, by decoupling the green and red lists splitting with the prior tokens, Unigram is twice as robust as KGW. But it’s more likely to leak the pattern of the watermarked tokens given that it uses a global green-red list splitting.
Exp. The Exp watermarking scheme from [KTHL23] is an extension of [Aar23]. Instead of using a single key as in KGW and Unigram, the usage of multiple watermark keys is inherent in Exp to provide the distortion-free guarantee. Each key is a vector of size $|\mathcal{V}|$ with values uniformly distributed in [0,1][0,1][ 0 , 1 ]. That is, $sk=\xi_{1},\xi_{2},\cdots,\xi_{n}$, where $\xi_{k}\in[0,1]^{|\mathcal{V}|},k\in[n]$, and $n$ is the length of the watermark keys, default to 256256256.
For the prediction of the token $\textbf{x}_{t+1}$, Exp firstly collects the output probability vector $\textbf{p}\in[0,1]^{|\mathcal{V}|}$ from the LLM. A random shift $r\overset{{\scriptscriptstyle\$}}{\leftarrow}[n]$ is sampled at the beginning of receiving the prompt. Then the token $\textbf{x}_{t+1}$ is sampled using the Gumbel trick [Gum48]:
@@
where $k=r+t+1\text{ mod }n$, i.e., each position uses a different watermark key which determines the uniform distribution sampling used in the Gumbel trick sampling. This method guarantees that the output distribution is distortion-free, whose expectation is identical to the distribution without a watermark given sufficiently large $n$.
@@
where $k=t\text{ mod }n$. Exp computes the minimum Levenshtein distance using the basic test statistic as a cost (see Sec. 2.4 in [KTHL23]).
Instead of using single keys as KGW and Unigram, Exp uses multiple keys and incorporates the Gumbel trick to rigorously provide the distortion-free (unbiased) guarantee, whose expected output distribution over the key space is identical to the unwatermarked distribution.
Sentence Quality. Perplexity (PPL) is one of the most common metrics for evaluating language models. It can also be utilized to measure the quality of the sentences [ZALW24, KGW+23a] based on the oracle of high-quality language models. Formally, PPL returns the following quality score for an input sentence x:
@@
In our evaluation, we utilize the GPT3 [OWJ+22] as the oracle model to evaluate sentence quality.
Watermark Setups and Hyper-parameters. For KGW [KGW+23a] and Unigram [ZALW24] watermarks, we utilize the default parameters in [ZALW24], where the watermark strength is $\delta=2$, and the green list portion is $\gamma=0.5$. We employ a threshold of $T=4$ for these two watermarks. For the Exp watermark (referred to as Exp-edit in [KTHL23]), we use the default parameters, where the watermark key length is $n=256$ and the block size $k$ defaults to be identical to the token length. We set the p-value threshold for Exp to 0.050.050.05 in our experiments.
Appendix B Proof of Theorem 1
In the following, we prove the bound on the maximum number of tokens that are allowed to be inserted into a watermarked sentence for Unigram [ZALW24].

### v3

@@ 6.3 Defending Detection with Differential Privacy
In light of the issues above, we propose an effective defense using ideas from differential privacy (DP) [8] to counteract detection API based spoofing attacks. DP adds random noise to function results evaluated on private dataset such that the results from neighbouring datasets are indistinguishable. Similarly, we consider adding Gaussian noise to the distance score in the watermark detection, making the detection $(\epsilon,\delta)$-DP [8], and ensuring that attackers cannot tell the difference between two queries by replacing a single token in the content, thus increasing the hardness of launching the attacks. Considering an attacker can average multiple query results to reduce noise and estimate original scores without DP protection, we propose to calculate the noise based on the random seed generated by a pseudorandom function (PRF) with the sentence to be detected as the input. Specifically, $\mathtt{seed}=\mathtt{PRF}_{sk}(\textbf{x})$, where $sk$ is the secret key held by the detection service. The users without the secret key cannot reverse or reduce the noise in the detection score. Thus, we can successfully mitigate the noise reduction via averaging multiple query results without comprising on utility or protection of the DP defense. In the following, we evaluate the utility of the DP defense and its performance in mitigating the spoofing attacks.
(d) Spoofing ASR and detection ACC.
(e) Z-scores with/without DP.
Figure 4: Evaluation of DP detection on KGW watermark and LLAMA-2-7B model. (a). Spoofing attack success rate (ASR) and detection accuracy (ACC) without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP. We use the best $\sigma=4$ from (a).
Experiment Setup. Firstly, we assess the utility of DP defense by evaluating the accuracy of the detection under various noise scales. Next, we evaluate the efficacy of the spoofing against DP detection defense using the same method as in Sec. 6.1. We select the optimal noise scale that provides best defense while keeping the drop in accuracy within $2\%$. We note that for KGW and Unigram watermarks, we add noise to the z-scores. Sensitivity varies with sentence length (e.g., $\Delta=\frac{h+1}{\sqrt{\gamma(1-\gamma)l}}$ for replacement editing, where $l$ is the sentence length, $h$ is the context width of the watermark, and $\gamma$ is the portion of the tokens in green list). The actual noise scale is proportional to $\sigma\Delta$.
Evaluation Result. As shown in Fig. 4, with a noise scale of $\sigma=4$, the DP detection’s accuracy drops from the original $98.2\%$ to $97.2\%$ on KGW and LLAMA-2-7B, while the spoofing ASR becomes $0\%$ using the same attack procedure as Sec. 6.1. The results are consistent for Unigram and Exp and OPT-1.3B as shown in Appendix L, which illustrates that the DP defense has a great utility-defense trade-off, with a negligible accuracy drop and significantly mitigates the spoofing.
@@ 6.4 Discussion
Guideline #3 Public detection APIs can enable both spoofing and removal attacks. To defend against these attacks, we propose a DP-inspired defense, which combined with techniques such as anomaly detection, query rate limiting, and user identification verification can help to make public detection more feasible in practice.
@@
The detection API, available to the public, aids users in differentiating between AI and human-created materials. However, it can be exploited by attackers to gradually remove watermarks or launch spoofing attacks. We propose a defense utilizing the ideas in differential privacy, which significantly increases the difficulty for spoofing attacks. However, this method is less effective against watermark-removal attacks that exploit the detection API because attackers’ actions will be close to random sampling, which, even though with less success rates, remains an effective way of removing watermarks. Therefore, we leave developing a more powerful defense mechanism against watermark-removal attacks exploiting detection API as future work. Additionally, we note that the attacker may increase the sensitivity of the input sentences by substituting multiple tokens and infer whether these tokens are in the green list or not to launch the spoofing attack, but this will require much more queries to the detection API. We recommend detection services adopt a defense-in-depth approach [2, 9]. E.g., they should detect and curb malicious behavior by limiting query rates from potential attackers, and also verify the identity of the users to protect against Sybil attacks.
7 Discussion, Limitation & Future Work
Generalizability of our attacks. We focus on three SOTA PRF-based robust watermarks, which are a natural set to explore given their popularity and formal guarantees. There are other watermarks like the semantics-based watermarks [24, 34]. While attacking semantics-based watermarks is outside the scope of our study, we deem this an interesting future direction to explore.
Recently, researchers have also proposed signature-based publicly detectable watermarks [10] to mitigate the spoofing attacks by exploiting robustness. Unlike the watermarks we study, these watermarks usually have weaker robustness guarantees, which further highlights the trade-offs between robustness and vulnerability to spoofing attacks, as we have discussed in Sec. 4. A more recent work [5] proposed a new watermarking scheme to embed undetectable watermarks. To achieve the undetectability, instead of relying on using multiple watermark keys as the popular watermarks we study [17, 46, 20], they use pseudorandom error-correcting codes, which are computationally indistinguishable from uniform random strings but can be detected using the trapdoor matrix. Given that they do not require using multiple watermark keys to mitigate watermark stealing, our removal attacks exploiting the use of multiple keys cannot be directly generalized to this watermarking scheme. We deem it an interesting direction to explore attack opportunities on such watermarking schemes.
Our findings, such as exploiting robustness properties and publicly available detection APIs, can also be generalized to image watermarks [40, 43]. The attackers must integrate domain-specific constraints to ensure that the generated sentences or images are meaningful and high-quality. We deem studying the fundamental trade-offs for image watermarks a promising future direction.
Trade-offs of watermark context width. There are two effective strategies to mitigate the watermark stealing attacks for the KGW watermark [17]: 1) using a larger context width $h$ and 2) using multiple watermark keys. In this work (Sec. 5), we primarily explore the fundamental trade-offs in using multiple watermark keys, which prior work has underexplored. Trade-offs in context widths were discussed in previous work [15, 17, 46]. Using a larger $h$ increases resistance against watermark stealing but reduces robustness. Recent work [15] shows successful watermark stealing even with $h=4$. Using multiple keys, as shown in Sec. 5 of our paper, mitigates stealing attacks but introduces new attack vectors of watermark removal. Our attacks will work under different choices of context width, as we exploit properties or design choices orthogonal to the context width. To demonstrate this point, we provide more experimental results in Appendix M.
The influence of how detection proceeds with multiple keys. For the scenarios where multiple keys are used, we consider the detector using min/max aggregation to obtain the detection score. More robust aggregations exist including the Harmonic mean p-value [41]. We note that our watermark-removal attack exploiting the use of multiple keys is not dependent on the aggregation method as we do not rely on the server’s watermark detection. However, the trade-off analysis and the sweet spot for the number of keys may slightly change given the different detection performance.
Changing to p-values in KGW and Unigram. P-values are used for Exp [20] watermark in our paper, and the observations are consistent with KGW [17] and Unigram [46]. We expect no impact on results from this change since p-values are monotonic to z-scores. To ease the figures’ presentation, we adopt the z-statistics in the main paper, we present more results of using p-values in Appendix M.
8 Conclusion
@@
In this work, we reveal new attack vectors that exploit common features and design choices of LLM watermarks. In particular, while these design choices may enhance robustness, resistance against watermark stealing attacks, and public detection ease, they also allow malicious actors to launch attacks that can easily remove the watermark or damage the model’s reputation. Based on the theoretical and empirical analysis of our attacks, we suggest guidelines for designing and deploying LLM watermarks along with possible defenses to establish more reliable LLM watermark systems.
Acknowledgements
This work was supported in part by the National Science Foundation grants IIS2145670, CCF2107024, CNS2326312 and funding from Amazon, Apple, Google, Intel, Meta, and the CyLab Security and Privacy Institute. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies.
Appendix A Broader Impacts
@@
Our work studies the security implications of common LLM watermarking design choices. By developing realistic attacks and defenses and a simple set of guidelines for watermarking in practice, we aim for the work to serve as a resource for the development of secure LLM watermarking systems. Of course, by outlining such attacks, there is a risk that our work may in fact increase the prevalence of watermark removal or spoofing attacks performed in practice. We believe that this is nonetheless an important step towards educating the community about potential risks in watermarking systems and ultimately creating more effective defenses for secure LLM watermarking.
@@
More generally, our work shows that a number of trade-offs exist in LLM watermarking (e.g., between utility, usability, robustness, and susceptibility to removal or spoofing attacks). The guidelines we propose provide rough proposals for considering these trade-offs, but we note that how to best navigate each trade-off will depend on the application at hand. Considering strategies to best navigate this space for specific LLM watermarking applications is an important direction of future study.
Appendix B Attacker’s Motivation
We present two practical scenarios to motivate watermark-removal attacks: (i) A student or a journalist uses high-quality watermarked LLMs to write articles, but wants to remove the watermark to claim originality. (ii) A malicious company offering LLM services for clients, instead of developing their own LLMs, simply queries a watermarked LLM from a victim company and removes the watermark, potentially infringing upon IP rights of the victim company.
In piggyback and spoofing attacks, an attacker can damage the reputation of a victim company offering an LLM service. For example: (i) The attacker can use a spoofing attack to generate fake news or incorrect facts and post them on social media. By claiming the material is generated by the LLM from the benign company, the attacker can damage the reputation of the company and their model. (ii) The attacker can use the spoofing attack to inject malicious code into some public software. The code has the benign company’s watermark embedded, and the benign company may thus be at fault and have to bear responsibility for the actions.
Appendix C Watermarking Schemes & Hyper-Parameters
@@
In this section, we introduce the three watermarking schemes we evaluate in the paper—KGW [17], Unigram [46], and Exp [20]. We also introduce the perplexity, a metric to evaluate the sentence quality.
@@
KGW. In the KGW watermarking scheme, when generating the current token $\textbf{x}_{t+1}$, all the tokens in the vocabulary is pseudorandomly shuffled and split into two lists—the green list and the red list. The random seed used to determine the green and red lists is computed by a watermark secret key $sk$ and the prior $h$ tokens $\textbf{x}_{t-h-1}||\cdots||\textbf{x}_{t}$ using pseudorandom functions (PRFs):
$\textsc{seed}=F_{sk}(\textbf{x}_{t-h-1}||\cdots||\textbf{x}_{t}),$
@@
where $h$ is the context width of the watermark. We note that the choice of $h$ has minor influence on our attacks or defenses, as our algorithms are not dependent on $h$. Here we use their original algorithm with $h=1$. Then, the seed is used to split the vocabulary into the green and red lists of tokens, with $\gamma$ portion of tokens in the green list:
@@
Then, KGW generates a binary watermark mask vector for the current token prediction, which has the same size as the vocabulary. All the tokens in the green list $L_{\text{green}}$ have value $1$ in the mask, and all the tokens in the red list have value 00 in the mask:
@@
To embed the watermark, KGW add a constant to the logits of the LLM’s prediction for token $\textbf{x}_{t+1}$:
@@
where $g$ is the number of tokens in the green list, $l$ is the total number of tokens in the input token sequence, and $\gamma$ is the portion of the vocabulary tokens in the green list. Similar to the watermark embedding, the green and red lists for each token position are determined by watermark secret key and the token prior to the current token in the input token sequence.
Unigram. Similar to KGW, Unigram also splits the vocabulary into green and red lists and prioritize the tokens in the green list by adding a constant to the logits before computing the softmax. The difference is that Unigram uses global red and green lists instead of computing the green and red lists for each token. That is, the seed to shuffle the list is only determined by the watermark secret key and generated by a Pseudo-Random Generator (PRG):
@@
The watermark embedding and detection procedures are the same as KGW: Unigram first compute the watermark mask:
@@
where $g$ is the number of tokens in the green list, $l$ is the total number of tokens in the input token sequence, and $\gamma$ is the portion of the vocabulary tokens in the green list. According to the analysis in [46] and also consistent with our results in Sec. 4.1, by decoupling the green and red lists splitting with the prior tokens, Unigram is twice as robust as KGW. But it’s more likely to leak the pattern of the watermarked tokens given that it uses a global green-red list splitting.
Exp. The Exp watermarking scheme from [20] is an extension of [1]. Instead of using a single key as in KGW and Unigram, the usage of multiple watermark keys is inherent in Exp to provide the distortion-free guarantee. Each key is a vector of size $|\mathcal{V}|$ with values uniformly distributed in $[0,1]$. That is, $sk=\xi_{1},\xi_{2},\cdots,\xi_{n}$, where $\xi_{k}\in[0,1]^{|\mathcal{V}|},k\in[n]$, and $n$ is the length of the watermark keys, default to $256$.
For the prediction of the token $\textbf{x}_{t+1}$, Exp firstly collects the output probability vector $\textbf{p}\in[0,1]^{|\mathcal{V}|}$ from the LLM. A random shift $r\overset{{\scriptscriptstyle\$}}{\leftarrow}[n]$ is sampled at the beginning of receiving the prompt. Then the token $\textbf{x}_{t+1}$ is sampled using the Gumbel trick [12]:
@@
where $k=r+t+1\text{ mod }n$, i.e., each position uses a different watermark key which determines the uniform distribution sampling used in the Gumbel trick sampling. This method guarantees that the output distribution is distortion-free, whose expectation is identical to the distribution without watermark given sufficiently large $n$.
@@
where $k=t\text{ mod }n$. And Exp computes the minimum Levenshtein distance using the basic test statistic as a cost (see Sec. 2.4 in [20]).
Instead of using single keys as KGW and Unigram, Exp uses multiple keys and incorporates Gumbel trick to rigorously provide distortion-free (unbiased) guarantee, whose expected output distribution over the key space is identical to the unwatermarked distribution.
Sentence Quality. Perplexity (PPL) is one of the most common metrics for evaluating language models. It can also be utilized to measure the quality of the sentences [46, 17] based on the oracle of high-quality language models. Formally, PPL returns the following quality score for an input sentence x:
@@
In our evaluation, we utilize the GPT3 [32] as the oracle model to evaluate sentence quality.
Setups and Hyper-Parameters. For KGW [17] and Unigram [46] watermarks, we utilize the default parameters in [46], where the watermark strength is $\delta=2$, and the green list portion is $\gamma=0.5$. We employ a threshold of $T=4$ for these two watermarks with a single watermark key. For the scenarios where multiple keys are used, we calculate the thresholds to guarantee that the false positive rates (FPRs) are below 1e-3. For the Exp watermark (refered to as Exp-edit in [20]), we use the default parameters, where the watermark key length is $n=256$ and the block size $k$ is default to be identical to the token length. We set the p-value threshold for Exp to $0.05$ in our experiments.
For the spoofing attack exploiting detection APIs, we obtain the first three tokens with the highest probabilities from the unwatermarked LLM, and for the removal attack exploiting detection APIs, we obtain the first five tokens with the highest probabilities from the watermarked LLM. For watermark removal attacks exploiting detection APIs on KGW and Unigram, we increase the probabilities of the tokens that have the smallest detection confidence, and then sample from the modified probability distribution. For watermark removal attacks exploiting detection APIs on Exp, we simply sample the token that has the maximum p-value, but we will skip the tokens that have low probabilities (lower than $0.15$) when the detection p-value is high (higher than $0.1$). The different setup for the Exp watermark is required to ensure that we can produce high-quality sentences. For watermark spoofing attacks that exploit detection APIs, we sample the token that has the highest detection confidence for KGW, Unigram, and Exp watermarks.
We conduct the experiments on a cluster with 8 NVIDIA A100 GPUs, AMD EPYC 7763 64-Core CPU, and 1TB memory.
Appendix D Attack Feasibility Analysis of Piggyback Spoofing Exploiting Robustness
We study the bound on the maximum number of tokens that are allowed to be inserted or edited in a watermarked sentence, and we present the following theorem on Unigram watermark [46] due to its clean robustness guarantee:
Theorem 1 (Maximum insertion portion).
Consider a watermarked token sequence x of length $l$. The Unigram watermark z-score threshold is $T$, the portion of the tokens in the green list is $\gamma$, the detection z-score of x is $z$, and the number of inserted tokens is $s$. Then, to guarantee the expected z-score of the edited text is greater than $T$, it suffices to guarantee $\frac{s}{l}\leq\frac{z^{2}-T^{2}}{T^{2}}$.